a)Solve the following:

i) 2x/2-16=1/32       ii) log8(x-2)+log8(x)=1

b) The demand curve for a swivel chair is given by p=4000(3-q) dollars per swivel chair, where p is the price and q is the quantity, in thousands of swivel chair, demanded at that price. What quantity will be demanded if the price per swivel chair is $256.60?

 


1
Expert's answer
2021-11-02T16:55:35-0400

(a)


(i)

"2 ^{\frac {x} {2} -16} = \frac {1} {32} \\"


"2 ^{\frac {x} {2} -16} = \frac {1} {2^5} \\"


"2 ^{\frac {x} {2} -16} = 2^ {-5} \\
Equating \ the \ indices; \\
\frac{x} {2} - 16 = -5 \\
Grouping \ like \ terms \ (i.e \ adding \ 16 \ to \ both \ sides) \\
\frac{x} {2} - 16 + 16 = -5 + 16 \\
\frac{x} {2} = 11 \\ 
Multiplying \ both \ sides \ by \ 2 \\
\frac{x} {2} \times 2 = 11 \times 2 \\
x = 22"


(ii)

"\log_8 (x -2) + \log_8 (x) = 1 \\
According \ to \ the \ rule \ of \ logarithm, \\
\log A + \log B = \log (A \times B) \\
\therefore \log_8 (x -2) + \log_8 (x) = 1 \\ \log_8 x(x -2) = 1 \\
\log_8 x(x -2) = \log_8 8 \\
Equating\ the \ logarithm, \\
x(x -2) = 8 \\
x^2 - 2x = 8 \\
x^2 - 2x - 8 = 0 \\
Solving \ using \ the \ quadratic \ formula; \\
x = \frac {-b \pm \sqrt {b^2 - 4ac}} {2a} \\
a = 1, b = -2, c = -8 \\
x = \frac {-(-2) \pm \sqrt {(-2)^2 - 4(1)(-8)}} {2 \times 1} \\
x = \frac {2 \pm \sqrt {4 + 32}} {2} \\
x = \frac {2 \pm \sqrt {36}} {2} \\
x = \frac {2 \pm 6} {2} \\
x = \frac {2 + 6} {2} \ or \ x = \frac {2 - 6} {2} \\
x = \frac {8} {2} \ or \ x = \frac {-4} {2} \\
x = 4 \ or \ x = -2"

since logarithm of negative number is not possible, (i.e., "x \isin \mathbb{N}")

"\therefore x = 4"


(b)

"p = 4000(3^{-q}) \\
\frac {p} {4000} = 3^{-q}"


Taking the logarithm to base 3 of both sides;

"\log_3 \frac {p} {4000} = \log_3 3^{-q} \\
\log_3 \frac {p} {4000} = -q \log_3 3 \\
since \ \log_3 3 =1 \\
\log_3 \frac {p} {4000} = -q \\
So, \\
q = - \log_3 \frac {p} {4000} \\

\therefore"

The quantity that will be demanded if the price per swivel chair is $256.60 is calculated by:


"q = - \log_3 \frac {256.60} {4000}"


changing logarithm of base three to base ten;

"\log_a b = \frac {\log_c (b)} {\log_c (a)}"


"q = - \log_3 \frac {256.60} {4000} = - \frac {\log_{10} \frac {256.60} {4000}} {\log_{10} 3}"


"q = - \frac {(-1.1928)} {0.4771}"


"q = 2.5001 \\
q \approx 2.5"


Therefore, 2.5 thousand swivel chairs will be demanded with the price of $256.60